In lewis Carrol's "Through the Looking Glass," Tweedledum says to Tweedledee, "The sum of your weight and twice mine is 361 pounds." Then Tweedledee says to Tweedledum, "Contrariwise, the sum of your weight and twice mine is 362 pounds." Find the weight of Tweedledum and Tweedledee.
Found 2 solutions by Earlsdon, Abbey:
Answer by Earlsdon(6294)

Did they really say this in that story?
Let x = Tweedledee's weight.
Let y = Tweedledum's weight.
So, according to Tweedledum, x + 2y = 361 Rewrite as: x = 361 - 2y and substitute in the equation below:
and according to Tweedledee, y + 2x = 362
After the substitution, we have:
y + 2(361 - 2y) = 362 Now, solving for y:
y + 722 - 4y = 362
-3y = -360
y = 120 lbs This is Tweedledum's weight, which is not too bad in lbs, but if it were in stones, he would definitely be obese.
x = 361 - 2y
x = 361 - 2(120)
x = 361 - 240
x = 121 lbs This is Tweedledee's weight and the same comment applies.
Check:
Tweedledum says, "The sum of your weight and twice mine is 361"
Since x is Tweedledee's weight, then we have:
x + 2y = 121 + 2(120) = 121 + 240 = 361 and, contrariwise, according to Tweedledee:
y + 2x = 120 + 2(121) = 120 + 242 = 362

Answer by Abbey(339)

let x - tweedledums weight
y= tweedledee's weight
x+2y=361
2x+y=362
Multiply the first equation by -2
-2x-4Y=-722
2x+y=362
Add them:
-3y=-360
divide by -3
y-120
x+2(120)=361
x+(240)=361
x=361-240
x=121
Check:
2*121+120=362
2*120+121=361
tweedledee weighs 120 and tweedledum weighs 121